40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 15 10:52:38 2019, gautam, Update, SUS, EY door opened 
    Reply  Tue Jan 15 15:49:47 2019, gautam, Update, SUS, EY door opened EY_OSEMs.png
       Reply  Sat Jan 19 15:33:18 2019, gautam, Update, SUS, ETMY OSEMs faulty Screen_Shot_2019-01-19_at_3.32.35_PM.pngETMY_sensors_1231832635.pdf
          Reply  Tue Jan 22 20:36:53 2019, gautam, Update, SUS, ETMY OSEMs faulty 
             Reply  Wed Jan 23 12:39:18 2019, gautam, Update, SUS, EY chamber work 6x
                Reply  Wed Jan 23 23:12:44 2019, gautam, Update, SUS, Prep for FC cleaning IY_level_before.pdf
                   Reply  Thu Jan 24 15:32:31 2019, gautam, Update, SUS, Y arm cavity side first contact applied Yarm_FC.pdfOSEMfilter.png
                      Reply  Tue Jan 29 22:12:40 2019, gautam, Update, SUS, Alignment prep 
                         Reply  Wed Jan 30 11:54:24 2019, gautam, Update, SUS, More alignment prep IMG_5958.JPGIMG_5962.JPG
                            Reply  Wed Jan 30 19:25:40 2019, gautam, Update, SUS, Xarm cavity alignment Xarm.pngXarm_IR.png
                               Reply  Fri Feb 1 01:24:06 2019, gautam, Update, SUS, Almost ready for pumpdown tomorrow 
                                  Reply  Fri Feb 1 13:16:50 2019, gautam, Update, SUS, Pumpdown 83 underway 
                                     Reply  Fri Feb 1 21:52:57 2019, gautam, Update, SUS, Pumpdown 83 underway 
                                  Reply  Fri Feb 1 14:44:14 2019, gautam, Update, SUS, Y arm FC cleaning and reinstall IMG_5974.JPGIMG_5986.JPGIMG_5992.JPG
    Reply  Wed Jan 16 16:25:25 2019, gautam, Update, SUS, Yarm locked Yarm_locked.png
       Reply  Fri Jan 18 21:34:18 2019, gautam, Update, SUS, Unused optic on EY table IMG_5408.JPG
          Reply  Sat Jan 19 05:07:45 2019, Koji, Update, SUS, Unused optic on EY table 
Message ID: 14423     Entry time: Wed Jan 30 11:54:24 2019     In reply to: 14422     Reply to this: 14424
Author: gautam 
Type: Update 
Category: SUS 
Subject: More alignment prep 

[chub, gautam]

  1. ETMY cage was wiped down
    • Targeted potential areas where dust could drift off from and get attracted to a charged HR surface
    • These areas were surprisingly dusty, even left a grey mark on the wipe [Attachment #1] - we think we did a sufficiently thorough job, but unclear if this helps the loss numbers
    • More pictures are on gPhoto
  2. Filters on SD and LR OSEMs were replaced - the open shadow sensor voltages with filters in/out are consistent with the T>95% coating spec.
  3. IPANG beam position was checked 
    • It is already too high, missing the first steering optic by ~0.5 inch, not the greatest photo but conclusion holds [Attachment #2].
    • I think we shouldn't worry about it for this pumpdown, we can fix it when we put in the new PR3.
  4. Cage wiping procedure was repeated on ITMY
    • The cage was much dustier than ETMY
    • However, the optic itself (barrel and edge of HR face) was cleaner
    • All accessible areas were wiped with isopropanol
    • Before/after pics are on gPhoto (even after cleaning, there are some marks on the suspension that looks like dust, but these are machining marks)

Procedure tomorrow [comments / suggestions welcome]:

  1. Start with IY chamber
    • Peel first contact with TopGun jet flowing
    • Inspect optic face with green flashlight to check for residual First Contact
    • Replace ITMY suspension cage in its position, clamp it down
    • Release ITMY from its EQ stops
    • Replace OSEMs in ITMY cage, best effort to recover previous alignment of OSEMs in their holders (I have a photo before removal of OSEMs), which supposedly minimized the coupling of the B-R modes into the shadow sensor signals
    • Best effort to have shadow sensor PD outputs at half their fully open voltages (with DC bias voltage applied)
    • Quick check that we are hitting the center of the ITM with the alignment tool
    • Check that the Oplev HeNe is reasonably centered on steering mirrors
    • Tie down OSEM cabling to the ITMY cage with clean copper wire
    • Replace the OSEM wiring tower
    • Release the SRM from its EQ stops
    • Check table leveling
    • Take pictures of everything, check that we have not left any tools inside the chamber
    • Heavy doors on
  2. Next, EY chamber
    • Repeat first seven bullets from the IY chamber, :%s/ITMY/ETMY/g
    • Confirm sufficient clearance between IFO beam axis and the elliptical reflector
    • Check Oplev beam path
    • Check table leveling
    • Take pictures of everything, check that we have not left any tools inside the chamber
    • Heavy doors on
  3. IFO alignment checks - basically follow the wiki, we want to be able to lock both arms (or at least see TEM00 resonances), and see that the PRC and SRC mode flashes look reasonable.
  4. Tighten all heavy doors up
  5. Pump down

All photos have been uploaded to google photos.

Attachment 1: IMG_5958.JPG  4.784 MB  Uploaded Wed Jan 30 18:47:03 2019  | Hide | Hide all
IMG_5958.JPG
Attachment 2: IMG_5962.JPG  5.911 MB  Uploaded Wed Jan 30 18:48:10 2019  | Hide | Hide all
IMG_5962.JPG
ELOG V3.1.3-